OpenStudy (anonymous):

what are the soulutions of (x+2) (x-3)=0

OpenStudy (anonymous):

x=-2 x=3

OpenStudy (anonymous):

(x+2)(x+3)=0 Either x+2=0 gives x=-2 Or x+3=0,gives x=-3 x=-2,-3

OpenStudy (anonymous):

(x+2)(x-3)=0 Either x+2=0,gives x=-2 Or x-3=0 ,gives x=3 solutions are x=-2,x=3